OpenStudy (anonymous):

What is the simplest form of the number:

OpenStudy (anonymous):

\[\sqrt{2}\left( \sqrt[8]{2} \right)\]

OpenStudy (anonymous):

Step by Step please

OpenStudy (anonymous):

the index is an 8 btw

OpenStudy (anonymous):

\[\sqrt{2}\sqrt[8]{2}=\sqrt[8]{2^4*2}=\sqrt[8]{2^5}\]

OpenStudy (anonymous):

http://awesomescreenshot.com/0132bkyjec

OpenStudy (anonymous):

5/8

OpenStudy (anonymous):

\[\sqrt{2}* \sqrt[8]{2}=2^{\frac{ 1 }{ 2 }}*2^{\frac{ 1 }{8 }}=2^{\left( \frac{ 1 }{ 2 }+\frac{ 1 }{ 8 } \right)}\] \[=2^{\frac{ 4+1 }{8 }}=2^{\frac{ 5 }{ 8 }}\] \[=\sqrt[8]{2^{5}}\]

OpenStudy (anonymous):

Ohhh understood... thank you both.

OpenStudy (anonymous):

yw